# The Complexity of Defining and Separating Fixpoint Formulae in Modal Logic

### v1

@@ 1. Introduction
For given logics $\mathcal{L},\mathcal{L}^{+}$, the $\mathcal{L}$-separability problem for $\mathcal{L}^{+}$ is to decide given two $\mathcal{L}^{+}$-formulae $\varphi,\varphi^{\prime}$ whether there is an $\mathcal{L}$-formula $\psi$ that separates $\varphi$ and $\varphi^{\prime}$ in the sense that $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. Obviously, a separator can only exist when $\varphi$ and $\varphi^{\prime}$ are mutually exclusive, and the problem is only meaningful when $\mathcal{L}$ is less expressive than $\mathcal{L}^{+}$. Intuitively, a separator formulated in a “simpler” logic $\mathcal{L}$ explains a given inconsistency in a “complicated” logic $\mathcal{L}^{+}$. Note that, for logics $\mathcal{L}^{+}$ closed under negation, $\mathcal{L}$-separability generalizes the $\mathcal{L}$-definability problem for $\mathcal{L}^{+}$: decide whether a given $\mathcal{L}^{+}$-formula is equivalent to an $\mathcal{L}$-formula. Indeed, $\varphi\in\mathcal{L}^{+}$ is equivalent to an $\mathcal{L}$-formula iff $\varphi$ and $\neg\varphi$ are $\mathcal{L}$-separable. Since separability is more general than definability, solving it requires an even better understanding of the logics under consideration. Both separability and definability are central problems with many applications in computer science. As seminal work let us only mention definability and separability of regular word languages by languages recognized by first-order logic [PZ16, Sch65, CH91].
In this paper we study definability and separability of formulae of the modal $\mu$-calculus $\mu\mathsf{ML}$ [Pra81, Koz83] by formulae in propositional modal logic $\mathsf{ML}$. $\mu\mathsf{ML}$ is the extension of $\mathsf{ML}$ with fixpoints that encompasses virtually all specification languages such as $\mathsf{PDL}$ [FL79] and $\mathsf{LTL}$ and $\mathsf{CTL}$ [BK08]. Let us consider an example.
@@
We explore the definability and separability problems over several classes of models relevant for computer science: all models, words, trees of bounded or unbounded outdegree; as well as restrictions of all these classes to finite models. On top of analyzing the decision problems, we also address the problem of constructing definitions and separators whenever they exist. The starting point for our research is the seminal paper of Otto [Ott99], where he solves modal definability over models of bounded and unbounded outdegree. In this paper, we continue this line of work, extend it to separability, and establish a fairly complete and interesting picture. Table 1 summarizes our results. We now explain its content further.
@@
$\mathsf{ML}$-def. ExpTime [Ott99] PSpace ExpTime [Ott99] ExpTime [Ott99]
@@
The first line essentially repeats Otto’s results; we only add the observation that $\mathsf{ML}$-definability over words is PSpace-complete. The second line contains our results for $\mathsf{ML}$-separability. The case of words is the easiest one, both in terms of computational complexity and required arguments. Next come the cases of binary and of unrestricted trees. These two classes possess some nice structural properties which (although true for different reasons) enable a common algorithmic treatment. Finally, the cases of trees with outdegree bounded by a number $d\geq 3$ enter the stage. These trees lack the good properties essential for previous constructions, which results in significantly higher computational complexity. The hardness result for $d\geq 3$ is interesting for two reasons. First, as it is entirely standard to encode trees of higher outdegree into binary ones, one could expect the ternary (and higher) case to have the same complexity as the binary one. And second, even though there are known cases when separation is provably harder than definability (regularity of visibly pushdown languages is decidable [LS19, Theorem 19] but regular separability thereof is not [Kop16, Theorem 2.4]), to the best of our knowledge our results are the only such case known in the area of logic.
@@
The complexity landscape for deciding separability is also reflected in the maximal sizes of the separators that we construct. Relying on the well-known connection of $\mu\mathsf{ML}$ to automata, we provide effective constructions for the cases of all models, words, and binary trees. It is worth mentioning that equally effective constructions for definability over all models are given in [LQ15], but they do not work for separability. The ternary case follows from a general argument. Our construction of separators over words is optimal. Under mild assumptions (there are at least two modalities) the constructions over binary and over unrestricted trees are optimal as well, but we leave it open whether these assumptions are needed for the lower bounds. In the case of ternary and higher outdegree trees we only conjecture optimality of the constructed separators.
@@
Finally, we observe that $\mathsf{ML}$ lacks the Craig interpolation property over trees of outdegree bounded by $d\geq 3$. Recall that a Craig interpolant for $\varphi\models\varphi^{\prime}$ in some logic $\mathcal{L}$ is a formula $\psi\in\mathcal{L}$ only using the common symbols of $\varphi$ and $\varphi^{\prime}$ and such that $\varphi\models\psi\models\varphi^{\prime}$. A logic satisfies the Craig interpolation property (CIP) if a Craig interpolant of $\varphi\models\varphi^{\prime}$ always exists. It is known that $\mathsf{ML}$ enjoys CIP over all models and over words [Gab72] and it follows from our techniques that this transfers to binary trees. In contrast and as mentioned above, over ternary and higher-arity trees $\mathsf{ML}$ lacks the CIP. It is worth mentioning that modal logic over frames of arity bounded by some $d$ has been studied under the name $\mathbf{K}\oplus\textit{{alt}}_{d}$ [Bel88]. Our results imply that $\mathbf{K}\oplus\textit{{alt}}_{d}$ enjoys CIP iff $d\leq 2$. Motivated by the lack of CIP over higher-arity trees, we study the induced interpolant existence problem – determining whether two given $\mathsf{ML}$-formulae $\varphi,\varphi^{\prime}$ admit a Craig interpolant – as a special case of separability. We show it to be coNExpTime-complete over higher arity trees, and thus harder than validity. Interpolant existence has recently been studied for other logics without CIP and our proofs are inspired by the techniques used there [JW21, AJM+23].
@@
As an application of our results for $d$-ary trees with $d\geq 3$ we additionally present a case study: definability and separability in the graded setting in which we allow modalities saying “there are at least $k$ children such that […]” [Fin72]. Such graded modalities are a standard extension of modal logic that is especially relevant in applications in knowledge representation for conceptual modeling [BHLS17]. Hence, the larger logic $\mathcal{L}^{+}$ is the graded $\mu$-calculus in this case, see for example [KSV02, CGL99]. For the separating language $\mathcal{L}$, we consider two options. In the first, we allow graded modalities also in the separating language graded modal logic $\mathsf{grML}$, whereas in the second we stick to $\mathsf{ML}$ as before. The first setting behaves essentially as the non-graded case in the sense that both $\mathsf{grML}$-definability and $\mathsf{grML}$-separability are ExpTime-complete. More interesting is the second option, where definability and separability again show different behaviour. While $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae is ExpTime-complete, $\mathsf{ML}$-separability is 2-ExpTime-complete. The intuitive reason for the higher complexity in the latter case is that trees of bounded arity are definable in graded $\mu\mathsf{ML}$. These results are also related with a recent study about separating logics supporting counting quantifiers by logics without these [KTWZ25].
@@
It is worth to mention that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae generalizes the boundedness problem which asks whether a formula with a single fixpoint is equivalent to a modal formula. Boundedness has been studied for other logics such as monadic-second order logic [BOW14], datalog [HKMV95], and the guarded fragment of first-order logic [BtCCV15]. Our paper is an extension of the preliminary papers [JK24, JK25].
@@ 2. Preliminaries
The semantics of $\mu\mathsf{ML}$ can be defined in multiple equivalent ways. The one most convenient for us is through parity games. These are infinite-duration games with perfect information, played between players $\exists\text{ve}$ and $\forall\text{dam}$. Parity games provide a convenient definition of the $\mu\mathsf{ML}$ semantics, and will be useful in some automata-theoretic constructions. We employ a standard terminology regarding parity games, such as positions, finite and infinite plays, strategies etc. (see [Ven20] for an introduction with a special focus on $\mu\mathsf{ML}$). Additionally, given a strategy $\zeta$ for either player we call a play $\pi$ a $\zeta$-play if $\pi$ is consistent with $\zeta$. Towards the game-theoretic definition of the semantics of $\mu\mathsf{ML}$ assume a model $\mathcal{M}$ and a formula $\varphi\in\mu\mathsf{ML}$. We define a parity game called the semantic game for $\mathcal{M}$ and $\varphi$ and denoted $\mathcal{G}(\mathcal{M},\varphi)$. The positions are $M\times\mathsf{sub}(\varphi)$. The moves depend on the topmost connective. From a position of the shape $(v,\psi\vee\psi^{\prime})$ or $(v,\psi\wedge\psi^{\prime})$ it is allowed to move to either $(v,\psi)$ or $(v,\psi^{\prime})$. From $(v,\Diamond\psi)$ and $(v,\Box\psi)$ the allowed moves lead to all $(w,\psi)$ such that $v\rightarrow w$. In position $(v,a)$ or $(v,\neg a)$ the game stops and $\exists\text{ve}$ wins iff $v$ satisfies the formula component $a$ or $\neg a$, respectively. From $(v,\mu x.\psi)$ and $(v,\mu x.\psi)$ the game moves to $(v,\psi)$, and from $(v,x)$ to $(v,\psi)$ where $\psi$ is the unique superformula of $x$ beginning with $\mu x$ or $\nu x$. $\exists\text{ve}$ owns positions whose formula component has $\vee$ or $\Diamond$ as the topmost connective and $\forall\text{dam}$ owns all other positions. $\exists\text{ve}$ wins an infinite play $\pi$ if the outermost subformula seen infinitely often in $\pi$ begins with $\nu$. We say that $\mathcal{M},v$ satisfies $\varphi$ and write $\mathcal{M},v\models\varphi$ if $\exists\text{ve}$ wins the game $\mathcal{G}(\mathcal{M},\varphi)$ from position $(v,\varphi)$. Since $\mathcal{M}$ is by definition pointed, we abbreviate $\mathcal{M},v_{I}\models\varphi$ with $\mathcal{M}\models\varphi$.
@@
In the paper we will study models of bounded and unbounded outdegree. The outdegree of a point $w\in M$ in a model $\mathcal{M}=(M,v_{I},\to,\mathsf{val})$ is the number of successors of $w$ in the underlying directed graph $G_{\mathcal{M}}=(M,\to)$. We say that $\mathcal{M}$ has bounded outdegree if there is a finite uniform upper bound $d$ on the outdegree of its points. In the latter case, we will call $\mathcal{M}$ $d$-ary, and binary or ternary if $d=2$ or $d=3$. If $d=1$, then we call $\mathcal{M}$ a word. A $d$-ary model is full if each of its nodes is either a leaf (i.e. has no children) or has precisely $d$ children. A model $\mathcal{M}$ is a tree if $G_{\mathcal{M}}$ is a (directed) tree with root $v_{I}$. We denote with $\mathbb{T}^{d}$ the class of all $d$-ary tree models. Both $\mathsf{ML}$ and $\mu\mathsf{ML}$ are invariant under bisimulation, and every ($d$-ary) model is bisimilar to a ($d$-ary) tree. Hence, we do not loose generality by only looking at tree models.
@@ 2.2. Automata
We exploit the well-known connection of $\mu\mathsf{ML}$ and automata that read tree models. A nondeterministic parity tree automaton (NPTA) over $d$-ary trees $\mathbb{T}^{d}$ is a tuple $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ where $Q$ is a finite set of states, $q_{I}\in Q$ is the initial state, $\Sigma=\mathcal{P}(\sigma)$ for some signature $\sigma\subseteq\mathsf{Prop}$, $\mathsf{rank}$ assigns each state a priority, and $\delta$ is a transition function of type:
@@
Then, consistency of $\rho$ with $\delta$ means that $\{\rho(v^{\prime})\ |\ v^{\prime}\in V\}\in\delta(\rho(v),\mathsf{val}(v)\cap\sigma)$ for every $v$ with a set $V$ of children. In either case, we call the run $\rho$ accepting if for every infinite path $v_{0},v_{1}\ldots$ in $\mathcal{M}$ the sequence $\mathsf{rank}(\rho(v_{0})),\mathsf{rank}(\rho(v_{1})),\ldots$ satisfies the parity condition. We write $\mathcal{M}\models\mathcal{A}$ in case $\mathcal{A}$ has an accepting run on $\mathcal{M}$. An automaton that is identical to $\mathcal{A}$ except that the original initial state is replaced with $q$ is denoted $\mathcal{A}[q_{I}\mapsfrom q]$. The size of an automaton $\mathcal{A}$ is the number of its states and is denoted by $|\mathcal{A}|$.
An NPTA $\mathcal{A}$ is equivalent to a formula $\varphi\in\mu\mathsf{ML}$ over a class $\mathbf{C}$ of trees when $\mathcal{M}\models\varphi$ iff $\mathcal{M}\models\mathcal{A}$ for every tree $\mathcal{M}\in\mathbf{C}$. We rely on the following classical result (see for example the discussion in [Var98] and the well-presented Dealternation Theorem 5.7 in [BC18]):
@@
For every $\mu\mathsf{ML}$-formula $\varphi$ and class $\mathbf{C}$ of trees, we can construct an NPTA with exponentially many states equivalent to $\varphi$ over $\mathbf{C}$. The construction takes exponential time when $\mathbf{C}\subseteq\mathbb{T}^{d}$ for some $d$, and doubly exponential time in the unrestricted case.

### v2

@@ 1. Introduction
For given logics $\mathcal{L},\mathcal{L}^{+}$, the $\mathcal{L}$-separability problem for $\mathcal{L}^{+}$ is to decide, given two $\mathcal{L}^{+}$-formulae $\varphi,\varphi^{\prime}$, whether there is an $\mathcal{L}$-formula $\psi$ that separates $\varphi$ and $\varphi^{\prime}$ in the sense that $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. Obviously, a separator can only exist when $\varphi$ and $\varphi^{\prime}$ are mutually exclusive, and the problem is only meaningful when $\mathcal{L}$ is less expressive than $\mathcal{L}^{+}$. Intuitively, a separator formulated in a “simpler” logic $\mathcal{L}$ explains a given inconsistency in a “complicated” logic $\mathcal{L}^{+}$. Note that, for logics $\mathcal{L}^{+}$ closed under negation, $\mathcal{L}$-separability generalizes the $\mathcal{L}$-definability problem for $\mathcal{L}^{+}$: decide whether a given $\mathcal{L}^{+}$-formula is equivalent to an $\mathcal{L}$-formula. Indeed, $\varphi\in\mathcal{L}^{+}$ is equivalent to an $\mathcal{L}$-formula iff $\varphi$ and $\neg\varphi$ are $\mathcal{L}$-separable. Since separability is more general than definability, solving it requires an even better understanding of the logics under consideration. Both separability and definability are central problems with many applications in computer science. As seminal work let us only mention definability and separability of regular word languages by languages recognized by first-order logic [32, 35, 11].
In this paper we study definability and separability of formulae of the modal $\mu$-calculus $\mu\mathsf{ML}$ [33, 25] by formulae in propositional modal logic $\mathsf{ML}$. $\mu\mathsf{ML}$ is the extension of $\mathsf{ML}$ with fixpoints that encompasses virtually all specification languages such as $\mathsf{PDL}$ [15] and $\mathsf{LTL}$ and $\mathsf{CTL}$ [3]. Let us consider an example.
@@
We explore the definability and separability problems over several classes of models relevant for computer science: all models, words, trees of bounded or unbounded outdegree; as well as restrictions of all these classes to finite models. On top of analyzing the decision problems, we also address the problem of constructing definitions and separators whenever they exist. The starting point for our research is the seminal paper of Otto [30], where he solves modal definability over models of bounded and unbounded outdegree. In this paper, we continue this line of work, extend it to separability, and establish a fairly complete and interesting picture. Table 1 summarizes our results. We now explain its content further.
@@
$\mathsf{ML}$-def. ExpTime [30] PSpace ExpTime [30] ExpTime [30]
@@
The first line essentially repeats Otto’s results; we only add the observation that $\mathsf{ML}$-definability over words is PSpace-complete. The second line contains our results for $\mathsf{ML}$-separability. The case of words is the easiest one, both in terms of computational complexity and required arguments. Next come the cases of binary and of unrestricted trees. These two classes possess some nice structural properties which (although true for different reasons) enable a common algorithmic treatment. Finally, the cases of trees with outdegree bounded by a number $d\geq 3$ enter the stage. These trees lack the good properties essential for previous constructions, which results in significantly higher computational complexity. The hardness result for $d\geq 3$ is interesting for two reasons. First, as it is entirely standard to encode trees of higher outdegree into binary ones, one could expect the ternary (and higher) case to have the same complexity as the binary one. And second, even though there are known cases when separation is provably harder than definability (regularity of visibly pushdown languages is decidable [29, Theorem 19] but regular separability thereof is not [24, Theorem 2.4]), to the best of our knowledge our results are the only such case known in the area of logic.
@@
The complexity landscape for deciding separability is also reflected in the maximal sizes of the separators that we construct. Relying on the well-known connection of $\mu\mathsf{ML}$ to automata, we provide effective constructions for the cases of all models, words, and binary trees. It is worth mentioning that equally effective constructions for definability over all models are given in [28], but they do not work for separability. The ternary case follows from a general argument. Our construction of separators over words is optimal. Under mild assumptions (there are at least two modalities) the constructions over binary and over unrestricted trees are optimal as well, but we leave it open whether these assumptions are needed for the lower bounds. In the case of ternary and higher outdegree trees we only conjecture optimality of the constructed separators.
@@
Finally, we observe that $\mathsf{ML}$ lacks the Craig interpolation property over trees of outdegree bounded by $d\geq 3$. Recall that a Craig interpolant for $\varphi\models\varphi^{\prime}$ in some logic $\mathcal{L}$ is a formula $\psi\in\mathcal{L}$ only using the common symbols of $\varphi$ and $\varphi^{\prime}$ and such that $\varphi\models\psi\models\varphi^{\prime}$. A logic satisfies the Craig interpolation property (CIP) if a Craig interpolant of $\varphi\models\varphi^{\prime}$ always exists. It is known that $\mathsf{ML}$ enjoys CIP over all models and over words [18] and it follows from our techniques that this transfers to binary trees. In contrast and as mentioned above, over ternary and higher-arity trees $\mathsf{ML}$ lacks the CIP. It is worth mentioning that modal logic over frames of arity bounded by some $d$ has been studied under the name $\mathbf{K}\oplus\textit{{alt}}_{d}$ [4]. Our results imply that $\mathbf{K}\oplus\textit{{alt}}_{d}$ enjoys CIP iff $d\leq 2$. Motivated by the lack of CIP over higher-arity trees, we study the induced interpolant existence problem – determining whether two given $\mathsf{ML}$-formulae $\varphi,\varphi^{\prime}$ admit a Craig interpolant – as a special case of separability. We show it to be coNExpTime-complete over higher arity trees, and thus harder than validity. Interpolant existence has recently been studied for other logics without CIP and our proofs are inspired by the techniques used there [23, 1].
@@
As an application of our results for $d$-ary trees with $d\geq 3$ we additionally present a case study: definability and separability in the graded setting in which we allow modalities saying “there are at least $k$ children such that […]” [14]. Such graded modalities are a standard extension of modal logic that is especially relevant in applications in knowledge representation for conceptual modeling [2]. Hence, the larger logic $\mathcal{L}^{+}$ is the graded $\mu$-calculus in this case, see for example [27, 9]. For the separating language $\mathcal{L}$, we consider two options. In the first, we allow graded modalities also in the separating language graded modal logic $\mathsf{grML}$, whereas in the second we stick to $\mathsf{ML}$ as before. The first setting behaves essentially as the non-graded case in the sense that both $\mathsf{grML}$-definability and $\mathsf{grML}$-separability are ExpTime-complete. More interesting is the second option, where definability and separability again show different behaviour. While $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae is ExpTime-complete, $\mathsf{ML}$-separability is 2-ExpTime-complete. The intuitive reason for the higher complexity in the latter case is that trees of bounded arity are definable in graded $\mu\mathsf{ML}$. These results are also related with a recent study about separating logics supporting counting quantifiers by logics without these [26].
@@
It is worth to mention that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae generalizes the boundedness problem which asks whether a formula with a single fixpoint is equivalent to a modal formula. Boundedness has been studied for other logics such as monadic-second order logic [6], datalog [19], and the guarded fragment of first-order logic [5]. Our paper is an extension of the preliminary papers [21, 22].
@@ 2. Preliminaries
The semantics of $\mu\mathsf{ML}$ can be defined in multiple equivalent ways. The one most convenient for us is through parity games. These are infinite-duration games with perfect information, played between players $\exists\text{ve}$ and $\forall\text{dam}$. Parity games provide a convenient definition of the $\mu\mathsf{ML}$ semantics, and will be useful in some automata-theoretic constructions. We employ a standard terminology regarding parity games, such as positions, finite and infinite plays, strategies etc. (see [38] for an introduction with a special focus on $\mu\mathsf{ML}$). Additionally, given a strategy $\zeta$ for either player we call a play $\pi$ a $\zeta$-play if $\pi$ is consistent with $\zeta$. Towards the game-theoretic definition of the semantics of $\mu\mathsf{ML}$ assume a model $\mathcal{M}$ and a formula $\varphi\in\mu\mathsf{ML}$. We define a parity game called the semantic game for $\mathcal{M}$ and $\varphi$ and denoted $\mathcal{G}(\mathcal{M},\varphi)$. The positions are $M\times\mathsf{sub}(\varphi)$. The moves depend on the topmost connective. From a position of the shape $(v,\psi\vee\psi^{\prime})$ or $(v,\psi\wedge\psi^{\prime})$ it is allowed to move to either $(v,\psi)$ or $(v,\psi^{\prime})$. From $(v,\Diamond\psi)$ and $(v,\Box\psi)$ the allowed moves lead to all $(w,\psi)$ such that $v\rightarrow w$. In position $(v,a)$ or $(v,\neg a)$ the game stops and $\exists\text{ve}$ wins iff $v$ satisfies the formula component $a$ or $\neg a$, respectively. From $(v,\mu x.\psi)$ and $(v,\mu x.\psi)$ the game moves to $(v,\psi)$, and from $(v,x)$ to $(v,\psi)$ where $\psi$ is the unique superformula of $x$ beginning with $\mu x$ or $\nu x$. $\exists\text{ve}$ owns positions whose formula component has $\vee$ or $\Diamond$ as the topmost connective and $\forall\text{dam}$ owns all other positions. $\exists\text{ve}$ wins an infinite play $\pi$ if the outermost subformula seen infinitely often in $\pi$ begins with $\nu$. We say that $\mathcal{M},v$ satisfies $\varphi$ and write $\mathcal{M},v\models\varphi$ if $\exists\text{ve}$ wins the game $\mathcal{G}(\mathcal{M},\varphi)$ from position $(v,\varphi)$. Since $\mathcal{M}$ is by definition pointed, we abbreviate $\mathcal{M},v_{I}\models\varphi$ with $\mathcal{M}\models\varphi$.
@@
In the paper we will study models of bounded and unbounded outdegree. The outdegree of a point $w\in M$ in a model $\mathcal{M}=(M,v_{I},\to,\mathsf{val})$ is the number of successors of $w$ in the underlying directed graph $G_{\mathcal{M}}=(M,\to)$. We say that $\mathcal{M}$ has bounded outdegree if there is a finite uniform upper bound $d$ on the outdegree of its points. In the latter case, we will call $\mathcal{M}$ $d$-ary, and binary or ternary if $d=2$ or $d=3$. If $d=1$, then we call $\mathcal{M}$ a word. A $d$-ary model is full if each of its nodes is either a leaf (i.e. has no children) or has precisely $d$ children. A model $\mathcal{M}$ is a tree if $G_{\mathcal{M}}$ is a (directed) tree with root $v_{I}$. We denote with $\mathbb{T}^{d}$ the class of all $d$-ary tree models. Both $\mathsf{ML}$ and $\mu\mathsf{ML}$ are invariant under bisimulation, and every ($d$-ary) model is bisimilar to a ($d$-ary) tree. Hence, we do not lose generality by only looking at tree models.
@@ 2.2. Automata
We exploit the well-known connection of $\mu\mathsf{ML}$ and automata that read tree models. A nondeterministic parity tree automaton (NPTA) over $d$-ary trees $\mathbb{T}^{d}$ is a tuple $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ where $Q$ is a finite set of states, $q_{I}\in Q$ is the initial state, $\Sigma=\mathcal{P}(\sigma)$ for some signature $\sigma\subseteq\mathsf{Prop}$, $\mathsf{rank}$ assigns each state a priority from $\mathbb{N}$, and $\delta$ is a transition function of type:
@@
Then, consistency of $\rho$ with $\delta$ means that $\{\rho(v^{\prime})\ |\ v^{\prime}\in V\}\in\delta(\rho(v),\mathsf{val}(v)\cap\sigma)$ for every $v$ with a set $V$ of children. In either case, we call the run $\rho$ accepting if for every infinite path $v_{0},v_{1}\ldots$ in $\mathcal{M}$ the sequence $\mathsf{rank}(\rho(v_{0})),\mathsf{rank}(\rho(v_{1})),\ldots$ satisfies the parity condition, that is, the minimal number that occurs infinitely often in this sequence is even. We write $\mathcal{M}\models\mathcal{A}$ and say that $\mathcal{A}$ accepts $\mathcal{M}$ in case $\mathcal{A}$ has an accepting run on $\mathcal{M}$. The set of all trees accepted by $\mathcal{A}$ is denoted $\mathcal{L}(\mathcal{A})$. An automaton that is identical to $\mathcal{A}$ except that the original initial state is replaced with $q$ is denoted $\mathcal{A}[q_{I}\mapsfrom q]$. The size of an automaton $\mathcal{A}$ is the number of its states and is denoted by $|\mathcal{A}|$. We will often consider the following trash normal form for $\mathcal{A}$: there is a distinguished trash state $q_{\bot}$ such that for every state $q$: $\mathcal{A}[q_{I}\mapsfrom q]$ accepts some model iff $q\neq q_{\bot}$.
An NPTA $\mathcal{A}$ is equivalent to a formula $\varphi\in\mu\mathsf{ML}$ over a class $\mathbf{C}$ of trees when $\mathcal{M}\models\varphi$ iff $\mathcal{M}\models\mathcal{A}$ for every tree $\mathcal{M}\in\mathbf{C}$. We rely on the following classical result:
@@
For every $\mu\mathsf{ML}$-formula $\varphi$ and class $\mathbf{C}$ of trees we can construct an NPTA $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ with alphabet $\Sigma=\mathcal{P}(\text{sig}(\varphi))$, size $|Q|$ exponential in $|\varphi|$, and number of ranks $|\{\mathsf{rank}(q)\mid q\in Q\}|$ polynomial in $|\varphi|$, such that $\mathcal{A}$ is equivalent to $\varphi$ over $\mathbf{C}$. The construction takes exponential time when $\mathbf{C}\subseteq\mathbb{T}^{d}$ for some $d$, and doubly exponential time in the unrestricted case. Moreover, in either case $\mathcal{A}$ is in trash normal form.
The construction of an automaton $\mathcal{A}$ equivalent to a given $\varphi$ with the claimed properties is entirely standard, see for example [37, Theorem 7] and the well-presented Dealternation Theorem 5.7 in [7]. Bringing $\mathcal{A}$ to trash normal form boils down to checking emptiness for every state. This is done in time exponential in $|\varphi|$ (or double exponential in the unrestricted case) thanks to the key observation underlying the single exponential algorithm for satisfiability of $\mu\mathsf{ML}$ [13]: although $\mathcal{A}$ has exponentially many states, its acceptance condition is simple. More specifically, in our terms this simplicity means that $\mathcal{A}$ has only polynomially many ranks. Emptiness reduces to a parity game with the same set of ranks and a number of positions polynomial (exponential, respectively) in $|Q|$. Since a parity game with $n$ positions and $m$ ranks can be solved in time $n^{O(m)}$ (in fact, even $n^{O(\log(m))}$, see [8, Main Theorem 2.7]), this gives us the desired complexity.